# Scalable Feature Learning on Huge Knowledge Graphs for Downstream Machine Learning

### v1

@@ 3 SEPAL: expanding from a core subgraph
SEPAL’s key idea is to allocate more computation time to the more frequent entities and then use message passing to propagate embeddings at low cost to regions of the graph where they have not been computed yet. It departs from existing embedding propagation methods [Vashishth et al., 2020, Wang et al., 2022] that compute embeddings on the full graph and use propagation as post-processing to smooth them. SEPAL is compatible with any embedding model whose scoring function has the form given by Equation 1, examples of which are provided in Table 1.
@@ 3.1 Splitting large graphs into manageable subgraphs
SEPAL first defines the core of a knowledge graph. The quality of the core embeddings is particularly important, as they serve as (fixed) boundary conditions during the propagation phase. To optimize this quality, two key factors must be considered during core selection: 1) ensuring a dense core subgraph by selecting the most central entities and 2) achieving full coverage of relation types. Yet, there can be a trade-off between these two objectives, hence, SEPAL offers two core selection procedures:
Degree-based selection: This simple approach selects the top entities by degree –with proportion $\eta_{n}\in(0,1)$– and keeps only the largest connected component of the induced subgraph. The resulting core is dense, which boosts performance for entity-centric tasks like feature enrichment (Section F.2). However, it does not necessarily contain all the relation types.
Hybrid selection: To ensure full relational coverage, this method combines two sampling strategies. First, it selects entities with the highest $\eta_{n}$ degrees. Second, for each relation type, it includes entities involved in edges with the highest $\eta_{e}$ degrees (where degree is the sum of the head and tail nodes’ degrees). The union of these two sets forms the core, and if disconnected, SEPAL reconnects it by adding the necessary entities (details in Section F.2). Hyperparameters $\eta_{n},\eta_{e}\in(0,1)$ are proportions of nodes and edges that control the core size.
Compared to degree-based selection, hybrid selection benefits tasks relying on relation embeddings, such as link prediction. However, its additional relation-specific edge sampling and reconnection steps can be computationally expensive for knowledge graphs with many relations.
@@
Extracting such subgraphs is challenging on large knowledge graphs. These are scale-free graphs with millions of nodes and no well-defined clusters [Leskovec et al., 2009]. They pose difficulties to partitioning algorithms. For instance, algorithms based on propagation, eigenvalues, or power iterations of the adjacency matrix [Raghavan et al., 2007, Shi and Malik, 2000, Newman, 2006] struggle with the presence of extremely high-degree nodes that make the adjacency matrix ill-conditioned. None of the existing partitioning algorithms satisfy our full set of constraints, and thus we devise our own algorithm, called BLOCS and described in detail in Appendix E. To satisfy the requirements despite these challenges, BLOCS creates overlapping subgraphs.
@@ 3.3 Outside the core: relation-aware propagation
where $\mathcal{N}(u)$ denotes the set of neighbors of outer entity $u$, $\mathcal{K}$ the set of positive triples of the graph, and $\alpha$ a hyperparameter similar to a learning rate. During updates, $\ell_{2}$ normalization projects embeddings on the unit sphere. With DistMult, this accelerates convergence by canceling the effect of neighbors that still have zero embeddings. Normalizing embeddings is a common practice in knowledge-graph embedding [Bordes et al., 2013, Yang et al., 2015], and SEPAL acts accordingly. During propagation, the core embeddings remain frozen.
4 Theoretical analysis
@@
Proposition 4.1 shows that SEPAL with DistMult minimizes an energy that only accounts for the positive triples. The more aligned the embeddings within positive triples, the lower this energy. In self-supervised learning, negative sampling is needed to prevent embeddings from collapsing to a single point [Hafidi et al., 2022]. However, in our case, this oversmoothing is avoided thanks to the boundary conditions of fixed core-entities and relations embeddings, which act as “anchors” in the embedding space.
@@
For an outer entity $u$, the gradient update of its embedding $\boldsymbol{\theta}_{u}^{(t+1)}=\boldsymbol{\theta}_{u}^{(t)}-\eta\frac{\partial\mathcal{E}}{\partial\boldsymbol{\theta}_{u}^{(t)}}$ boils down to SEPAL’s propagation equation, for a learning rate $\eta=\alpha$. See Section D.2 for detailed derivations. ∎
@@
Section D.1 shows that SEPAL’s propagation, when combined with DistMult, resembles an Arnoldi-like iteration, suggesting that entity embeddings converge toward generalized eigenvectors of an operator that integrates both global graph structure and relation semantics (via fixed relation embeddings).
@@
Proposition 4.1 shows that SEPAL minimizes an energy that promotes global alignment between embeddings of positive triples, leading to embeddings with high queriability. In contrast, classic KGE methods use negative sampling to incorporate a supplementary constraint of local contrast between positive and negative triples. This adds discriminative power to the model, useful for link prediction, but can be detrimental to the global alignment of embeddings, which is important for queriability.
5 Experimental study: utility to downstream task and scalability
@@
To compare large knowledge graphs of different sizes, we use Freebase [Bollacker et al., 2008], WikiKG90Mv2 [(an extract of Wikidata) Hu et al., 2020], and three generations of YAGO: YAGO3 [Mahdisoltani et al., 2014], YAGO4 [Pellissier Tanon et al., 2020], and YAGO4.5 [Suchanek et al., 2024]. We expand YAGO4 and YAGO4.5 into a larger version containing also the taxonomy, i.e., types and classes –which algorithms will treat as entities– and their relations. We discard numerical attributes and keep only the largest connected component (Section A.1). To perform an ablation study of SEPAL without BLOCS for which we need smaller datasets, we also introduce Mini YAGO3, a subset of YAGO3 built with the 5% most frequent entities. Knowledge graph sizes are reported in Figure 5.
@@
We evaluate the embeddings as node features in downstream tasks [Grover and Leskovec, 2016, Cvetkov-Iliev et al., 2023, Robinson et al., 2025, Ruiz et al., 2024]. Specifically, we incorporate the embeddings in tables as extra features and measure the prediction improvement of a standard estimator in regression tasks. This setup allows us to compare the utility of knowledge graphs of varying sizes. Indeed, for a task, a suboptimal embedding of a larger knowledge graph may be more interesting than a high-quality embedding of a smaller knowledge graph because the larger graph brings richer information, on more entities. We benchmark 4 downstream regression tasks [adapted from Cvetkov-Iliev et al., 2023]: Movie revenues, US accidents, US elections, and housing prices. Details are provided in Section A.2.
Figure 2: Statistical performance on real-world tables. a) Pareto frontiers of averaged normalized prediction scores with respect to embedding times (log-scale). b) Critical difference diagrams [Terpilowski, 2019] of average ranks among the three methods (SEPAL, PBG and DGL-KE) that scale to every knowledge-graph dataset. The ranks are averaged over all tasks; a task being defined as the combination of a downstream table and a source knowledge graph. SEPAL gets the best average downstream performance for each of the 7 source knowledge graphs. Figure 5 gives the detailed results for each table. Appendix B.1 details the metric used.
Larger knowledge graphs do bring value (Figure 2), as they cover more entities of the downstream tasks (Figure 4). For each source graph, SEPAL provides the best embeddings and is much more scalable (details in Figure 5). Considering performance/cost Pareto optimality across methods and source graphs (Figure 2a), SEPAL achieves the best performance with reduced cost, but the simple baseline FastRP also gives Pareto-optimal results, for smaller costs. Although FastRP discards the type of relation, it performs better than most dedicated KGE methods. Its iterations also solve a more global problem, like SEPAL (Section D.1).
We used DistMult as base model as it is a classic and good performer [Ruffinelli et al., 2020, Kadlec et al., 2017, Jain et al., 2020]. Figure 22 shows that SEPAL can speed up DistMult over 20 times for a given training configuration. For other scoring functions like RotatE and TransE, Figure 7 shows that SEPAL also improves the downstream performance of its base model.
@@
We also evaluate SEPAL on tables from WikiDBs [Vogel et al., 2024], a corpus of databases extracted from Wikidata. We build 42 downstream tables (26 classification tasks, and 16 regression tasks), described in Section A.2. Four of them are used as validation tasks, to tune hyperparameters, and the remaining 38 are used as test tables. Figure 3 presents the aggregated results on these 38 test tables, showing that, here also, applying SEPAL to very large graphs provides the best embeddings for downstream node regression and classification, and that SEPAL brings a decreased computational cost. Regarding the performance-cost tradeoff, FastRP is once again Pareto-optimal for small computation times, highlighting the benefits of global methods.
Figure 3: Statistical performance on WikiDBs tables. a) Critical difference diagrams of scalable methods. Black lines connect methods that are not significantly different. b) Pareto frontiers of averaged normalized prediction scores with respect to embedding times (log-scale). Figure 6 in Section C.2 provides the detailed results for each of the 38 test tables.
@@ 6 Discussion and conclusion
Our work focuses on embeddings for feature enrichment of downstream tables, an active research field [Cvetkov-Iliev et al., 2023, Ruiz et al., 2024, Robinson et al., 2025]. Another popular use case for embeddings is knowledge graph completion. However, this task is fundamentally different from feature enrichment: embeddings optimized for link prediction may not perform well for feature enrichment, and vice versa [Ruffinelli and Gemulla, 2024]. Nevertheless, we also evaluate SEPAL on knowledge graph completion, for which we expect lower performances given that SEPAL does not locally optimize the contrast between positive and negative triples scores. Results reported in Section C.4 show that SEPAL sometimes performs lower than existing methods (DGL-KE, PBG) on this task, although no method is consistently better than the others for all datasets, and statistical tests show no significant differences (Figure 8).
Conclusion: for downstream machine learning, embeddings should be optimized globally
@@
In this paper, we show how to optimize knowledge-graph embeddings for downstream machine learning. We propose a very scalable method, SEPAL, and conduct a comprehensive evaluation on 7 knowledge graphs and 46 downstream tables showing that SEPAL both: (1) markedly improves predictive performance on downstream tasks and (2) brings computational-performance benefits –multiple-fold decreased train times and bounded memory usage– when embedding large-scale knowledge graphs. Our theoretical analysis suggests that SEPAL’s strong performance on downstream tasks stems from its global optimization approach, resulting in better-aligned embeddings compared to classic methods based on negative sampling. SEPAL improves the quality of the generated node features when used for data enrichment in external (downstream) tasks, a setting that can strongly benefit from pre-training embeddings on knowledge bases as large as possible. It achieves this without requiring heavy engineering, such as distributed computing, and can easily be adapted to most KGE models.

### v2

@@ 3 SEPAL: revisiting knowledge-graph embedding optimization
SEPAL’s key idea is to allocate more computation time to the more frequent entities and then use message passing to propagate embeddings at low cost to regions of the graph where they have not been computed yet. It departs from existing embedding propagation methods (vashishth2020compositionbased; wang2022rep) that compute embeddings on the full graph and use propagation as post-processing to smooth them. SEPAL is compatible with any embedding model whose scoring function has the form given by Equation 2, examples of which are provided in Table 1.
@@ 3.1 Splitting large graphs into manageable subgraphs
SEPAL first defines the core of a knowledge graph. The quality of the core embeddings is particularly important, as they serve as (fixed) boundary conditions during the propagation phase. Good relation embeddings are also key to structuring the propagation. To optimize this quality, two key factors must be considered during core selection: 1) ensuring a dense core subgraph by selecting the most central entities and 2) achieving full coverage of relation types. Yet, there can be a trade-off between these two objectives, hence, SEPAL offers two core selection procedures:
Degree-based selection: This simple approach selects the top entities by degree –with proportion $\eta_{n}\in(0,1)$– and keeps only the largest connected component of the induced subgraph. The resulting core is dense, which boosts performance for entity-centric tasks like feature enrichment (Section˜F.2.2). However, it does not necessarily contain all the relation types.
Hybrid selection: To ensure full relational coverage, this method combines two sampling strategies. First, it selects entities with the highest $\eta_{n}$ degrees. Second, for each relation type, it includes entities involved in edges with the highest $\eta_{e}$ degrees (where degree is the sum of the head and tail nodes’ degrees). The union of these two sets forms the core, and if disconnected, SEPAL reconnects it by adding the necessary entities (details in Section˜F.2.2). Hyperparameters $\eta_{n},\eta_{e}\in(0,1)$ are proportions of nodes and edges that control the core size.
Compared to degree-based selection, hybrid selection benefits tasks relying on relation embeddings, such as link prediction. However, its additional relation-specific edge sampling and reconnection steps can be computationally expensive for knowledge graphs with many relations. For disconnected input graphs, all connected components other than the largest one are added to the core subgraph.
@@
Extracting such subgraphs is challenging on large knowledge graphs. These are scale-free graphs with millions of nodes and no well-defined clusters (leskovec2009community). They pose difficulties to partitioning algorithms. For instance, algorithms based on propagation, eigenvalues, or power iterations of the adjacency matrix (raghavan2007near; shi2000normalized; newman2006finding) struggle with the presence of extremely high-degree nodes that make the adjacency matrix ill-conditioned. None of the existing partitioning algorithms satisfy our full set of constraints, and thus we devise our own algorithm, called BLOCS and described in detail in Appendix˜E. To satisfy the requirements despite these challenges, BLOCS creates overlapping subgraphs.
BLOCS uses three base mechanisms to grow the subgraphs: diffuse (add all neighboring entities to the current subgraph), merge (merge two overlapping subgraphs), and dilate (add all unassigned neighboring entities to the current subgraph). There are two different regimes during the generation of subgraphs. First, few entities are assigned, and the computationally effective diffusion quickly covers a large part of the graph, especially entities that are close to high-degree nodes. However, once these close entities have been assigned, the effectiveness of diffusion drops because it struggles to reach entities farther away. For this reason, BLOCS switches from diffusion to dilation once the proportion of assigned entities reaches a certain threshold $h$ (a hyperparameter chosen $\approx.6$, depending on the dataset). By adding only unassigned neighbors to subgraphs, dilation drives subgraph growth towards unassigned distant entities. However, the presence of long chains can drastically slow down this regime because they make it add entities one by one. Some knowledge graphs have long chains, for instance YAGO4.5 (see Diameter in Table 2). To tackle them, BLOCS switches back to diffusion for a few steps, with seeds taken inside the long chains.
BLOCS works faster on graphs that have small diameters, where most entities can be reached during the diffusion regime and fewer dilation steps are required (Section˜F.1).
@@ 3.3 Outside the core: relation-aware propagation
where $\mathcal{N}(u)$ denotes the set of neighbors of outer entity $u$, $\mathcal{K}$ the set of positive triples of the graph, and $\alpha$ a hyperparameter similar to a learning rate. During updates, $\ell_{2}$ normalization projects embeddings on the unit sphere. With DistMult, this accelerates convergence by canceling the effect of neighbors that still have zero embeddings. Normalizing embeddings is a common practice in knowledge-graph embedding (bordes2013translating; yang2015embedding), and SEPAL acts accordingly. During propagation, the core embeddings remain frozen.
4 Theoretical analysis: embedding alignment
@@
Proposition˜4.1 shows that SEPAL with DistMult minimizes an energy that only accounts for the positive triples. The more aligned the embeddings within positive triples, the lower this energy. In self-supervised learning, negative sampling is needed to prevent embeddings from collapsing to a single point (hafidi2022negative). However, in our case, this oversmoothing is avoided thanks to the boundary conditions of fixed core-entities and relations embeddings, which act as “anchors” in the embedding space.
@@
As a consequence, SEPAL converges towards a stationary point of $\mathcal{E}$ on the unit sphere.
@@
For an outer entity $u$, the gradient update of its embedding $\boldsymbol{\theta}_{u}^{(t+1)}=\boldsymbol{\theta}_{u}^{(t)}-\eta\frac{\partial\mathcal{E}}{\partial\boldsymbol{\theta}_{u}^{(t)}}$ boils down to SEPAL’s propagation equation, for a learning rate $\eta=\alpha$. See Section˜D.2 for detailed derivations. ∎
@@
Section˜D.1 shows that SEPAL’s propagation, when combined with DistMult, resembles an Arnoldi-like iteration, suggesting that entity embeddings converge toward generalized eigenvectors of an operator that integrates both global graph structure and relation semantics (via fixed relation embeddings).
@@
Proposition˜4.1 shows that SEPAL minimizes an energy that promotes global alignment between embeddings of positive triples, which we argue leads to embeddings with high queriability. In contrast, classic KGE methods use negative sampling to incorporate a supplementary constraint of local contrast between positive and negative triples. This adds discriminative power to the model, useful for link prediction, but can be detrimental to the global alignment of embeddings, which is important for queriability.
5 Experimental study: utility to downstream tasks
@@
To compare large knowledge graphs of different sizes, we use Freebase (bollacker2008freebase), WikiKG90Mv2 ((an extract of Wikidata) hu2020open), and three generations of YAGO: YAGO3 (mahdisoltani2014yago3), YAGO4 (pellissier2020yago), and YAGO4.5 (suchanek2024yago). We expand YAGO4 and YAGO4.5 into a larger version containing also the taxonomy, i.e., types and classes –which algorithms will treat as entities– and their relations. We discard numerical attributes and keep only the largest connected component (Section˜A.1). To perform an ablation study of SEPAL without BLOCS for which we need smaller datasets, we also introduce Mini YAGO3, a subset of YAGO3 built with the 5% most frequent entities. Knowledge graph sizes are reported in Figure 5.
@@
We evaluate the embeddings as node features in downstream tasks (grover2016node2vec; cvetkov2023relational; robinson2025relbench; ruiz2024high). Specifically, we incorporate the embeddings in tables as extra features and measure the prediction improvement of a standard estimator in regression tasks. This setup allows us to compare the utility of knowledge graphs of varying sizes. Indeed, for a task, a suboptimal embedding of a larger knowledge graph may be more interesting than a high-quality embedding of a smaller knowledge graph because the larger graph brings richer information, on more entities. We benchmark 4 downstream regression tasks (adapted from cvetkov2023relational): Movie revenues, US accidents, US elections, and housing prices. Details are provided in Section˜A.2.
Figure 2: Statistical performance on real-world tables. a) Pareto frontiers of averaged normalized prediction scores with respect to embedding times (log-scale). b) Critical difference diagrams (terpilowski2019scikit) of average ranks among the three methods (SEPAL, PBG and DGL-KE) that scale to every knowledge-graph dataset. The ranks are averaged over all tasks; a task being defined as the combination of a downstream table and a source knowledge graph. SEPAL gets the best average downstream performance for each of the 7 source knowledge graphs. Figure 5 gives the detailed results for each table. Appendix B.1 details the metric used.
Larger knowledge graphs do bring value (Figure 2), as they cover more entities of the downstream tasks (Figure 4). For each source graph, SEPAL provides the best embeddings and is much more scalable (details in Figure 5). Considering performance/cost Pareto optimality across methods and source graphs (Figure 2a), SEPAL achieves the best performance with reduced cost, but the simple baseline FastRP also gives Pareto-optimal results, for smaller costs. Although FastRP discards the type of relation, it performs better than most dedicated KGE methods. Its iterations also solve a more global problem, like SEPAL (Section˜D.1).
We used DistMult as base model as it is a classic and good performer (Ruffinelli2020You; Kadlec2017KnowledgeBC; jain2020knowledge). Figure 22 shows that SEPAL can speed up DistMult over 20 times for a given training configuration. For other scoring functions like RotatE and TransE, Figure 7 shows that SEPAL also improves the downstream performance of its base model.
@@
We also evaluate SEPAL on tables from WikiDBs (vogel2024wikidbs), a corpus of databases extracted from Wikidata. We build 42 downstream tables (26 classification tasks, and 16 regression tasks), described in Section˜A.2. Four of them are used as validation tasks, to tune hyperparameters, and the remaining 38 are used as test tables. Figure 3 presents the aggregated results on these 38 test tables, showing that, here also, applying SEPAL to very large graphs provides the best embeddings for downstream node regression and classification, and that SEPAL brings a decreased computational cost. Regarding the performance-cost tradeoff, FastRP is once again Pareto-optimal for small computation times, highlighting the benefits of global methods.
Figure 3: Statistical performance on WikiDBs tables. a) Critical difference diagrams of scalable methods. Black lines connect methods that are not significantly different. b) Pareto frontiers of averaged normalized prediction scores with respect to embedding times (log-scale). Figure 6 in Section˜C.2 provides the detailed results for each of the 38 test tables.
@@ 6 Discussion and conclusion
Our work focuses on embeddings for feature enrichment of downstream tables, an active research field (cvetkov2023relational; ruiz2024high; robinson2025relbench). Another popular use case for embeddings is knowledge graph completion. However, this task is fundamentally different from feature enrichment: embeddings optimized for link prediction may not perform well for feature enrichment, and vice versa (ruffinelli2024beyond). Nevertheless, we also evaluate SEPAL on knowledge graph completion, for which we expect lower performances given that SEPAL does not locally optimize the contrast between positive and negative triples scores. Results reported in Section˜C.6 show that SEPAL sometimes performs lower than existing methods (DGL-KE, PBG) on this task, although no method is consistently better than the others for all datasets, and statistical tests show no significant differences (Figure 8).
Conclusion: embeddings for downstream machine learning
@@
In this paper, we show how to optimize knowledge-graph embeddings for downstream machine learning. We propose a highly scalable method, SEPAL, and conduct a comprehensive evaluation on 7 knowledge graphs and 46 downstream tables showing that SEPAL both: (1) markedly improves predictive performance on downstream tasks and (2) brings computational-performance benefits –multiple-fold decreased train times and bounded memory usage– when embedding large-scale knowledge graphs. Our theoretical analysis suggests that SEPAL’s strong performance on downstream tasks stems from its global optimization approach, resulting in better-aligned embeddings compared to classic methods based on negative sampling. SEPAL improves the quality of the generated node features when used for data enrichment in external (downstream) tasks, a setting that can strongly benefit from pre-training embeddings on knowledge bases as large as possible. It achieves this without requiring heavy engineering, such as distributed computing, and can easily be adapted to most KGE models.